wondering if anyone knows how to work this.

Determine the total area under the standard
normal curve


A. find the area under the normal curve to
the left of z= -2 plus the area under the
normal curve to the right of z= 2

B. Find the area under the normal curve
to the left of z= -1.53 plus the area under
the normal cure to the right of
z= 2.53

No other information was provided for this question.
Answer #1

Ans:

We have to find:

Area under the normal curve to the left of z= -2 plus the area under the normal curve to the right of z= 2:

P(z<-2)+P(z>2)=0.02275+0.02275=0.0455

use normal distribution table or excel function normsdist(z)